A freight train left Tutuban station a 4 am, travelling 30 miles per hour. At 6 am, an express train travelling 50 miles per hour left the same station. When did the rtrain overtake the freight train?
Answer by checkley77(12844)

D=RT
D=30T FOR THE FREIGHT TRAIN.
D=50(T-2)
THE DISTANCES ARE THE SAME, THUS SET THESE 2 EQUATIONS EQUAL.
30T=50(T-2)
30T=50T-100
30T-50T=-100
-20T=-100
T=-100/-20
T=5 HOURS.
4:00+5:00=9:00 P.M. THEY WILL MEET.
PROOF:
30*5=50(5-2)
150=50*3
150=150